Blizkost (Perl 5 ¡==¿ Perl 6)

Duke Leto What is Blizkost?

Blizkost

(Perl 5 ¡==¿ Perl 6)

Jonathan ”Duke” Leto

Blizkost (Perl 5 ¡==¿ Perl 6)

Duke Leto What is Blizkost?

Flavors of Perl 6

• Rakudo - Parrot

• Pugs - Haskell

• SMOP - C+DSLs

• Elf - Ruby

• STD.pm - Larry’s Implementation in Perl 6 (gimme5)

• v6.pm - Source filter for Perl 5

• others ...

Duke Leto What is Blizkost?

Thanks

• Larry

• Eric Wilhelm

• Jerry Gay

• Patrick Michaud

• The Perl Foundation

• Everyone working on Parrot, Rakudo and Perl 6

Duke Leto What is Blizkost?

Resources

• http://rakudo.org

• http://parrot.org

• http://perlsphere.net

• #perl6 and #parrot on freenode • #pdx.pm on irc.perl.org